OpenStudy (anonymous):

could someone please check my work and see where i went wrong Integral x^3/ (x^2+a^2) dx Let u=x^2+a^2 where x^2=(u-a^2) du=2x dx where dx=du/2x

OpenStudy (anonymous):

Integral x^3/ (x^2+a^2) dx = Integral x^3/ u du/2x = Integral x^2/u * du/2 = 1/2 Integral (u-a^2)/u * du = 1/2 Integral (u/u - a^2/u) du = 1/2 Integral (1-a^2/u) du = 1/2 Integral 1du - 1/2 Integral a^2/u du = 1/2x - 1/2a^2 Integral 1/u du = 1/2x - 1/2a^2 In(u) + c = 1/2x - 1/2a^2 In(x^2+a^2) + c

ganeshie8 (ganeshie8):

check below highlighted line again : = 1/2 Integral (u-a^2)/u * du = 1/2 Integral (u/u - a^2/u) du = 1/2 Integral (1-a^2/u) du = 1/2 Integral 1du - 1/2 Integral a^2/u du `= 1/2x - 1/2a^2 Integral 1/u du = 1/2x - 1/2a^2 In(u) + c` = 1/2x - 1/2a^2 In(x^2+a^2) + c

ganeshie8 (ganeshie8):

whats the integral of \(1\) with respect to \(u\) ?

OpenStudy (anonymous):

do you mean integral 1 du?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

exactly! \(\large \int 1~ du = ?\)

OpenStudy (anonymous):

it is x +c

ganeshie8 (ganeshie8):

who is x ?

ganeshie8 (ganeshie8):

we're integrating with respect to \(u\) right ? why bother about x here ?

OpenStudy (anonymous):

sorry u +c

ganeshie8 (ganeshie8):

Right, then do you see the mistake

OpenStudy (anonymous):

thank you so much i was a little bit confused

OpenStudy (anonymous):

ive been solving a lot of integrals i guess its time for bed

ganeshie8 (ganeshie8):

everything else is correct :) good job !!

ganeshie8 (ganeshie8):

wolfram is your best friend for checking integrals : http://www.wolframalpha.com/input/?i=Integral+x%5E3%2F+%28x%5E2%2Ba%5E2%29+dx+

OpenStudy (anonymous):

Thank you for that @ganeshie8 I really appreciate it

OpenStudy (anonymous):

i gave one small question when i plug in x^2+a^2 into u it comes up with a slightly different answer

OpenStudy (anonymous):

have*

OpenStudy (anonymous):

=1/2(x^2+a^2) - 1/2a^2 In(x^2+a^2) + c

ganeshie8 (ganeshie8):

somehow this is same as the answer given by wolfram :) maybe differentiate ur final answer and see if you're getting back the integrand or not

ganeshie8 (ganeshie8):

did u get why they are both equal ? :)

ganeshie8 (ganeshie8):

notice that \(\frac{1}{2}a^2\) is just a constant so you can pack it into the arbitrary constant \(c\) : \(\frac{1}{2}a^2 + c = C_2\

ganeshie8 (ganeshie8):

* \(\frac{1}{2}a^2 + c = C_2 \)

OpenStudy (anonymous):

ok so the a^2 has been taken out

ganeshie8 (ganeshie8):

yes because \(\frac{1}{2}a^2 + c\) is as arbitrary a constant as \(c\) there is no need for two constant terms in the final answer

OpenStudy (anonymous):

if i had a million medals i would of given them to you :)

ganeshie8 (ganeshie8):

you can always pack the constant terms like this in the final answer

OpenStudy (anonymous):

I thank you again

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

ill write it down as a note